# COOCO - Common Objects Out-of-Context - Semantic Violation in Scenes: Investigating Multimodal Context in Referential Communication

### v1

@@ 5 Results
Figure 4: RefCLIPScore and Accuracy across experimental conditions, aggregated over all models.
@@
Figure 3 presents the RefCLIPScore results for each model at noise level 0 across all relatedness conditions. Models achieve the highest scores in the original and same target conditions, indicating better alignment with the reference when the target object is preserved or replaced by one from the same category. Interestingly, performance is higher in the same target condition than in the original. In contrast, performance declines in conditions where the target object differs from the original (high, medium, and low relatedness). Among these, models perform best in the high relatedness condition, with scores dropping as relatedness decreases. For models depicted in warm colours (the first four), the low relatedness condition yields performance that is similar to, or slightly higher than, the medium condition. Conversely, models in cool colours (XGen-MM-Phi3 and Kosmos-2) show a clear and consistent decline in performance with decreasing relatedness.
@@
Figure 4(a) presents RefCLIPScore values aggregated across all models and organized by relatedness categories, noise areas, and noise levels. As expected, models perform best in the zero-noise condition, followed closely by the context noise conditions (noise levels 0.5 and 1 applied outside the target). Similar conclusions can be drawn from the accuracy results (Figure 4(b)), where the highest accuracy ($\sim$0.8) is observed in the original and same target conditions under zero noise and context noise (0.5 and 1.0). These are followed by the high relatedness condition (accuracy $\sim$0.5), and the medium and low conditions (accuracy $\sim$0.4). Interestingly, in conditions where the target is not related to the scene (medium and low relatedness), the addition of context noise slightly improves performance on RefCLIPScore (Figure 4(a)), suggesting that target objects that deviate from the scene semantics are better predicted when misleading contextual cues are reduced.
@@
In contrast, applying noise directly to the target patch results in a clear performance degradation, proportional to the noise intensity: the higher the noise level, the greater the drop. When noise (at 0.5) is applied to the target or the entire image, accuracy in the original and same target conditions drops to  0.5, while all other conditions approach zero (Figure 4(b)). For RefCLIPScore, we find that at 0.5 noise level, for the high, medium, and especially low relatedness conditions, performance is more robust when noise is applied to the entire image rather than solely to the target patch. At a higher noise level (1.0), performance continues to decline. However, models perform best with target-specific noise in the original and same target conditions. This suggests that given a non-perturbed scene context, model predictions for noised target regions are for objects that would be expected given the scene composition. Comparing the 0.5 and 1.0 all-noise settings, we see a sharper drop for the original and same-target conditions, while the drop is smaller for the high, medium and especially the low relatedness conditions.
Figure 5: Semantic similarity between outputs and scene labels for the LLaVA-One-Vision 0.5B model. These results are divided by correctness (blue: correct, orange: incorrect), relatedness, noise area, and noise level.
@@
[…]
6 Output vs. Scene Similarity
We follow up our analysis of overall results with an in-depth focus on one of the models, LLaVA-OneVision-0.5B. As shown in Figure 3, this model performs competitively. Its relatively small size also makes it feasible to analyse in more depth.
We begin in this section with an analysis of the semantic similarity between the model’s predicted reference to the target (e.g. laptop), and the scene label for an image (e.g. office), to evaluate whether the generation was more strongly anchored to the overall scene or to the target object itself. The results, shown in Figure 5, are separated by accuracy, into correct (blue bars) and incorrect (orange bars) responses.
A general trend emerges among both correct and incorrect outputs, namely that when relatedness is medium or low, the model predicts targets which are less semantically related to the scene label. This indicates that the model’s predictions in these conditions are less anchored by the scene semantics. Crucially, when noise is applied to the target region, this trend disappears for incorrect outputs. Notably, across all relatedness conditions involving target noise, correct responses show higher similarity to the scene compared to the no-noise condition. This confirms an observation in Section 5, namely, that when the target is degraded by noise, the model relies more on global scene context. In successful cases, this shift results in outputs that are more semantically aligned with the scene.
@@
We further analyse these results using a Linear Mixed Effects (LME) analysis, with semantic similarity between predicted target and scene label as the dependent variable. We include noise level (modeled as a continuous variable, scaled and centred) and noise area (comparing target and context noise conditions to the ‘all’ noise condition) as predictors, with random intercepts for scene labels. [^3] We do not include a full random effects structure, as introducing random slopes resulted in poor convergence. For this analysis, we do not separate correct and incorrect outputs, focusing instead on the similarity of the model’s prediction to the scene in the two conditions representing the ‘extreme’ cases of target-scene congruence, namely, the original and the low-relatedness conditions.
Table 1 summarises the LME results. The most important observations are related to the interactions between noise level and noised region: for original images, as noise increases in both the target region and the context region, there is a marginal increase in the semantic similarity of the predicted target to the scene label. For images in the low relatedness condition, increasing noise in the context region results in a significant decrease in similarity between model predictions and scene label.
7 Attention Deployment Analysis
Maintaining our focus on the 0.5B version of LLaVA-OneVision Li et al. (2025), we now turn to an analysis of the model’s attention dynamics under different experimental conditions.
@@ 7 Attention Deployment Analysis
To analyze attention allocation across different layers $l$ of the vision encoder and COOCO samples $t$, we compute the ratio $r$ between the total normalized attention values assigned to the target ($a_{\mathrm{target}}$) and the context ($a_{\mathrm{context}}$): $r_{l}^{(t)}=\frac{a_{\mathrm{target},l}^{(t)}}{a_{\mathrm{context},l}^{(t)}}$. A value of $r$ closer to 1 indicates a higher allocation of attention to the target region, whereas a value closer to 0 suggests greater attention to the context. Since the context area is larger than the target, we do not expect $r$ to exceed 0.5. Instead, our focus is on analyzing its variations across different experimental conditions. We then compute the mean ratio value for each encoder layer by averaging across all samples: $\bar{r}_{l}=\frac{1}{T}\sum_{t=1}^{T}r_{l}^{(t)}$ where $T$ denotes the total number of instances. This allows us to assess how attention is distributed across different hierarchical levels of the encoder under varying experimental conditions.
Full details of how we extract attention values and compute layerwise attention ratios are in Appendix C.
@@
Figure 6 shows the attention allocation ratio across layers $\bar{r}_{l}$ for different semantic relatedness conditions at zero noise, separated by correct and incorrect responses according to the accuracy metric (Section 4.3). First, we observe that, for correct responses, the model generally applies more attention to the target compared to incorrect ones. Second, target-focused attention is primarily concentrated in the middle layers, peaking between layers 7 and 15. Third, the model’s attention to the target is clearly modulated by its semantic relatedness to the scene: more attention is allocated in the low-relatedness condition, followed by the medium-relatedness condition, and lowest in the high-relatedness condition. The least attention is allocated to the target in the original condition, while the same target condition receives slightly more. Interestingly, this pattern of increased attention to less semantically related targets also appears in incorrect responses. This suggests that even when the model fails to correctly name the object, it may still detect its incongruity with the surrounding scene, allocating more attention to targets that appear out of place.
@@
Figure 7: Layer-wise attention heatmaps showing how the model’s focus on the target varies with noise level (rows: none, moderate, high) and noise location (columns: full input, context only, target only). Warmer colors indicate stronger attention relative to the no-noise baseline (average attention ratio $\sim$0.18), while cooler colors indicate reduced attention.
@@
Figure 7 compares mean attention-to-target ratio $\bar{r}_{l}$ heatmaps for correct versus incorrect model responses across the three noise levels (0.0, 0.5, 1.0) and three perturbation conditions (All Input, Context only, Target only). The analysis shows that moderate noise on the target (e.g., noise level 0.5 in the All Input and Target Only conditions) tends to increase attention to the target in correct responses particularly in the middle layers (7-14), suggesting that the model compensates for partial degradation by focusing more on the target. However, when the noise becomes too severe, attention to the target declines, likely because the target information becomes less recoverable. In contrast, under the Context Only condition, higher noise levels consistently lead to increased attention on the target, indicating a reallocation of focus away from the noisy context toward the unaltered target.
Figure 8: Difference in mean attention-to-target ratio ($\Delta\bar{r}_{l}$) between correct and incorrect model predictions across layers, per noise level. Green shades indicate higher attention in correct responses compared to incorrect ones, while pink shades indicate the opposite. The diverging colour map is centered at zero and scaled symmetrically to ±0.25, the maximum observed delta.
Higher attention to the target is linked to correct responses under noisy conditions.
Figure 8 visualises the change in mean attention-to-target ratio $\Delta\bar{r}_{l}$ between correct and incorrect responses, across all experimental conditions. The heatmaps are predominantly green, indicating that when the model correctly recognizes the object, it tends to allocate more attention to the target token compared to when it misclassifies the object. This effect becomes more pronounced under higher noise conditions, particularly in the middle layers of the transformer (7-14). However, this observation is affected by both noise and target-scene relatedness: when a medium ($0.5$) level of noise is applied to the target or the whole image, correct object categorisation is marked by notably higher attention to the target, compared to incorrect cases.
@@ 8 Discussion
Several patterns emerge from our experimental findings. Firstly, the RefCLIPScore analysis reveals that models consistently exhibit superior performance when either the original target object or a semantically similar one is present. This may be due to the COCO images having been included in the training data of all models. However, the observation that scores are slightly higher for the "same target" condition compared to the "original" condition suggests that models benefit from slight semantic shifts within the same category, perhaps because categorising an object similar but not identical to the original introduces nuanced semantic cues that models leverage e.g., the generated same-category object is more prototypical than the actual object, making it easier to classify for the model. Moreover, the drop in performance corresponding to decreasing relatedness shows how sensitive models are to object semantics. The accuracy data aligns well with RefCLIPScore patterns, emphasising the robustness of object categorisation in zero- and context-noise conditions. However, target-specific noise substantially diminishes performance, emphasising the models’ reliance on clear object representations.
@@
The output versus scene similarity results for the LLaVA-One-Vision model indicate that as semantic relatedness diminishes, models tend to anchor outputs more to the semantics of the target rather than the broader scene. When target objects are corrupted by noise, correct responses shift towards a stronger reliance on scene context, implying an adaptive strategy that leverages global context to compensate for degraded local features.
@@
In attention analysis, clear differences emerge in attention allocation. Correct categorisation consistently involves higher attention on the target, particularly in middle layers (7–15), suggesting critical semantic processing occurs there, as already pointed out by previous multimodal interpretability studies on the LLaVA model family Zhang et al. (2025a, b); Golovanevsky et al. (2025). A compelling observation is the inverse relationship between semantic relatedness and attention allocation: less scene-related targets receive more focused attention. Furthermore, noise conditions modulate attention dynamics significantly. Moderate target-specific noise (0.5 level) appears to trigger a compensatory increase in target-focused attention, aiding correct recognition. Conversely, extreme noise (level 1.0) forces models to shift reliance entirely to context. This dynamic attention behaviour highlights an adaptive mechanism within the models, balancing local target analysis against global context integration depending on image clarity and semantic coherence.
Overall, these findings suggest that vision-language models dynamically modulate their semantic and attentional strategies, demonstrating adaptability that hinges heavily on target clarity and semantic coherence with the visual scene. This suggests that models become sensitive to scene semantics during training.

### v2

@@ 5 Results across models
Figure 4: RefCLIPScore and Accuracy across experimental conditions, aggregated over all models. When differences within a condition are minimal, the leftmost point corresponds to the higher value, followed by lower values to the right.
@@
Figure 3 reports RefCLIPScore results for each model at noise level 0 across all relatedness conditions. Models obtain the highest scores in the original and same target conditions, indicating stronger alignment with the reference when the target object is preserved or replaced by an object from the same category. Interestingly, performance is higher in the same target condition than in the original. A possible reason for this is that inpainting introduces subtle artefacts that make it easier for target objects to be picked out. Performance decreases when the target object differs from the original (high, medium, and low relatedness). Within these conditions, models perform best under high relatedness, with scores progressively declining as relatedness decreases. Importantly, the observed effects of object–scene relatedness are also supported by results on VISIONS, where no inpainting is applied (see § 6).
For models depicted in warm colours (the first four in Figure 3), the low relatedness condition yields performance that is similar to, or slightly higher than, the medium condition. Conversely, models in cool colours (XGen-MM-Phi3 and Kosmos-2) show a more consistent decline in performance with decreasing relatedness.
@@
Figure 4(a) reports RefCLIPScore aggregated across models by relatedness, noise area, and noise level. Overall, models achieve their best performance at context noise 0.5, followed by noise 1.0 and then the zero-noise condition, with the notable exception of the low-relatedness condition, where noise 1.0 yields the highest alignment. Consistent trends are observed in the accuracy results (Figure 4(b)): the original and same target conditions reach the highest accuracy ($\sim$0.8) across noise settings, followed by the high-relatedness condition ($\sim$0.5) and the medium and low-relatedness conditions ($\sim$0.4). While most relatedness conditions peak at context noise 0.5, the low-relatedness condition shows a monotonic improvement with increasing noise, indicating that maximal removal of contextual information is particularly beneficial when targets violate scene semantics and scene context becomes detrimental.
@@
In contrast, applying noise directly to the target region results in a clear performance degradation, proportional to the noise intensity: the higher the noise level, the greater the drop. When noise (at 0.5) is applied to the target or the entire image, accuracy in the original and same target conditions drops to  0.5, while all other conditions approach zero (Figure 4(b)). For RefCLIPScore, we find that at 0.5 noise level, for the high, medium, and especially low relatedness conditions, performance is more robust when noise is applied to the entire image rather than solely to the target. At a higher noise level (1.0), performance continues to decline. However, models perform best with target-specific noise in the original and same target conditions. This suggests that, given a non-perturbed scene context, model predictions for noised target regions are for objects that would be expected given the scene composition (see § 6.1). Comparing the 0.5 and 1.0 all-noise settings, we see a sharper drop for the original and same-target conditions, while the drop is smaller for the high, medium and especially the low relatedness conditions.
Mixed-Effects Analysis of RefCLIP Similarity
Focusing on the ‘extreme’ semantic conditions (original and low relatedness), we build a Linear Mixed-Effects (LME) model with RefCLIPScore as the dependent variable, noise level (scaled and centred) as a continuous predictor, and noise area as a categorical factor contrasting context and target noise against the all-noise baseline. The model includes random intercepts for the image identifier and the VL models. [^4] Since we are interested here in establishing the statistical reliability of the trends observed above across all models. Table 1 reports the LME results. Confirming our observations, baseline RefCLIPScore is higher for the original than for the low-relatedness images. Context noise increases similarity, while target noise decreases it. As expected, this contrast is stronger in the low-relatedness condition. Noise intensity shows a significant negative main effect, with interactions showing a less steep negative impact when noise is in the context or the target compared to full-image noise.
Original images Low relatedness
Predictor Coef. Std.Err. z Coef. Std.Err. z
Intercept 0.735 0.002 379.21 0.692 0.003 275.51
Noise area (context) 0.052 0.001 94.13 0.062 0.000 141.32
Noise area (target) -0.015 0.001 -27.34 -0.047 0.000 -107.65
Noise level (cont.) -0.034 0.000 -145.42 -0.034 0.000 -182.28
Noise area (context) × Noise level 0.032 0.001 60.75 0.036 0.000 86.79
Noise area (target) × Noise level 0.014 0.001 26.76 0.024 0.000 56.84
Group Var 0.041 0.069
Table 1: Linear mixed effects model: Noise area and noise level on RefCLIPScores, for original and low-relatedness images. denotes a significant effect at $p\leq.001$.
Figure 5: Semantic similarity between the model’s outputs and scene labels for LLaVA-OneVision-0.5B in COOCo (ours) and VISIONS Allegretti et al. (2025), shown across correctness (blue: correct; pink/red: incorrect), semantic-fit conditions (COOCo: high/low; VISIONS: congruent/incongruent), noise areas (target, context, all), and noise levels (0.5, 1.0). Dashed and solid horizontal lines indicate baseline similarity at zero noise for incorrect and correct predictions, respectively.
@@
Table 2: Linear mixed effects model: Noise area and noise level on scene-target semantic similarity, for original and low relatedness contexts. denotes a significant effect at $p\leq.01$. Conditional $R^{2}$ values: 0.52 (original) and 0.39 (low rel.)
6 In-depth analysis of LLaVA-OneVision
We follow up with further analysis of LLaVA-OneVision-0.5B, chosen because of its size, which makes in-depth exploration feasible and its competitive performance (cf. Figure 3; full accuracy scores for this model are in Appendix B, Tab 6 and 7). We report analyses on both COOCo and VISIONS Allegretti et al. (2025), rescaling images in the latter from the original (1920 $\times$ 1080 px) to match the resolution in COOCo (max. width of 640 px). We note that for VISIONS, naming differences between dataset labels and human-provided modal object names have negligible impact on accuracy (see Appendix A.3).
6.1 Output vs. scene similarity
We begin with an analysis of the text-based semantic similarity metric scores between the model’s predicted reference to the target (e.g. laptop), and the scene label for an image (e.g. office), to evaluate whether the generation was more strongly anchored to the overall scene or to the target object itself. Figure 5 reports how semantic similarity between the model’s outputs and the scene labels varies across correctness, noise conditions and intensity, and the relatedness distinctions defined in each dataset: high vs. low relatedness in COOCo, and congruent vs. incongruent scenes in VISIONS. [^5] Appendix B, Fig. 10 presents a complementary analysis, treating object-scene relatedness as continuous rather than categorical.
When the target region contains no noise (in the baseline condition with 0 noise, and in the context-noise setup), the same trend appears in both correct and incorrect outputs: in low-relatedness conditions, the model produces predictions that are less semantically aligned with the scene label than in the higher-relatedness conditions of each dataset. This indicates that, when the target object violates scene semantics, predictions are less anchored to the global scene and are instead more influenced by the unrelated object itself. Crucially, this pattern disappears when noise is applied to the target region. Incorrect predictions under target noise no longer exhibit clear differences between high and low semantic fit, suggesting that once the target’s appearance is degraded, object semantics cease to guide outputs. In contrast, correct predictions under target noise consistently shift towards the scene label: across all semantic-fit conditions, correct responses exhibit higher similarity to the scene than in the no-noise baseline (solid line). This aligns with the observation in § 5 that degrading the target prompts the model to compensate by relying more heavily on global scene context. When successful, the resulting predictions are more semantically aligned with the scene. Notably, only in the congruent case of the VISIONS dataset, incorrect predictions fall below their zero-noise baseline, suggesting that congruent objects, when not degraded, may anchor the answers to a higher semantic similarity with the scene.
When noise is applied to the context region, correct responses exhibit no meaningful deviation from the zero-noise baseline, indicating that when the target remains intact, the model continues to anchor its prediction primarily to the object itself rather than the broader scene. Incorrect responses, however, show a general decrease in similarity relative to the baseline, suggesting that context degradation harms scene-based anchoring when the model fails to identify the target correctly. Under all noise conditions, the effects differ across datasets: COOCo exhibits patterns comparable to those observed under target noise alone, whereas in VISIONS, the incorrect predictions in congruent trials maintain their baseline similarity to the scene, and those in incongruent trials even show an increase in similarity.
@@ 6 In-depth analysis of LLaVA-OneVision
We further analyse these results using LME on the COOCo data; the same analysis on VISIONS is reported in Appendix B, Tab 5. We treat semantic similarity (logit-transformed) between predicted target and scene label as the dependent variable. Noise level (modeled as a continuous variable, scaled and centred) and area (comparing target and context to the all noise condition) are predictors, with random intercepts for scene labels. [^6] We do not include a full random effects structure, as introducing random slopes resulted in poor convergence. We include both correct and incorrect outputs, analysing separately the two ‘extreme’ cases of target-scene congruence, namely, the original and the low-relatedness conditions. Table 2 summarises the LME results. The main effect of noise level is significant in the original, but not the low relatednesscondition. The target noise area is signficantly different from the all noise condition, but context is only significant in low-relatedness images. (All main effects are significant in VISIONS; cf. Tab. 5). More important are the interactions: In original images, increased noise in the target or context region results in slight but significant increases in semantic similarity of the predicted target to the scene label. In the low relatedness condition, increasing noise in the context region results in a significant decrease in similarity between model predictions and scene label.
6.2 Attention Deployment Analysis
@@
We now turn to an analysis of the model’s attention dynamics under different experimental conditions. In what follows, we use the term ‘attention’ to refer to the Transformer decoder’s attention to the image tokens when it generates the name of the object. To analyze attention allocation across different layers $l$ of the vision encoder and image samples $t$, we compute the ratio $r$ between the total normalized attention values assigned to the target ($a_{\mathrm{target}}$) and the context ($a_{\mathrm{context}}$): $r_{l}^{(t)}=\frac{a_{\mathrm{target},l}^{(t)}}{a_{\mathrm{context},l}^{(t)}}$. A value of $r$ higher than 1 indicates a higher allocation of attention to the target region, whereas a value closer to 0 suggests greater attention to the context. Since the context area is larger than the target, we do not expect $r$ to exceed 0.5. Instead, our focus is on analysing its variations across different experimental conditions. We then compute the mean ratio value for each encoder layer by averaging across all samples: $\bar{r}_{l}=\frac{1}{T}\sum_{t=1}^{T}r_{l}^{(t)}$ where $T$ denotes the total number of instances. This allows us to assess how attention is distributed across different hierarchical levels of the encoder under varying experimental conditions. Full details of how we extract attention values and compute layerwise attention ratios are in Appendix A.4.
@@
Figure 6 shows the attention allocation ratio across layers $\bar{r}_{l}$ for different semantic relatedness conditions at zero noise, separated by correct and incorrect responses according to the accuracy metric (§ 4.3). We note several trends that hold in both COOCo and VISIONS datasets. First, we observe that, for correct (resp. incorrect) responses, the model generally applies more attention to the target. Second, target-focused attention is primarily concentrated in the middle layers, peaking between layers 7 and 15. Third, the model’s attention to the target is clearly modulated by its semantic congruency with the scene: more attention is allocated in the low-relatedness condition, followed by the medium-relatedness condition, and least in the high-relatedness condition. The least attention is allocated to the target in the original condition, while the same target condition receives slightly more. Interestingly, this pattern of increased attention to less semantically related targets also appears in incorrect responses. This suggests that even when the model fails to correctly name the object, it may still detect its incongruity with the surrounding scene, allocating more attention to targets that appear out of place. While the above observations apply to both datasets, we also note that $\bar{r}_{l}$ is generally lower in VISIONS. This is unlikely to be due to cues introduced by inpainting, as the difference also holds in comparison to the original image condition in COOCo. We hypothesise that the lowered attention-to-target in VISIONS is either an artefact of the image compression performed for our experiments, or is related to general properties of the images (e.g. the relative size of target regions). Deeper comparison of the differences between the two datasets is left to future work.
@@
Figure 7: Layer-wise attention heatmaps showing how the model’s focus on the target varies with noise level (rows: none; moderate with $\lambda=0.5$; high with $\lambda=1.0$) and noise location (columns: all, context only, target only). Warmer colors indicate stronger attention relative to the no-noise baseline (average attention ratio $\sim$0.19), while cooler colors indicate reduced attention to the target. Target-scene relatedness levels range from low to original for COOCo, and in/congruent for VISIONS.
@@
Figure 7 compares mean attention‐to‐target ratio $\bar{r}_{l}$ heatmaps for correct versus incorrect model responses across the three noise levels (0.0, 0.5, 1.0) and three noise conditions (all, context, target). The analysis shows that moderate noise on the target (e.g., noise level 0.5 in the all and target noise conditions) tends to increase attention to the target in correct responses, particularly in the middle layers (7-14), suggesting that the model compensates for partial degradation by focusing more on the target. However, when the noise becomes too severe, attention to the target declines, likely because the target information becomes less recoverable. In contrast, under the context noise condition, higher noise levels consistently lead to increased attention on the target, indicating a reallocation of focus away from the noisy context toward the unaltered target. These attention patterns appear consistently in both COOCo and VISIONS. Once again we note a subtle difference between the two datasets: in the all noise condition, there is virtually no attention to the target region in the VISIONS samples. One possible reason, already alluded to above, is that inpainting in COOCo creates subtle cues that facilitate attention to the target region. However, in the COOCo original (non-inpainted) images we do observe some attention to the target, albeit much less so than in other conditions. These observations suggest that artefacts due to inpainting may facilitate attention to the target under all noise, but they are not the sole determining factor.
Finally, we also see that when the model generates a correct response, there is higher attention to the target region, also under noisy conditions. Visualisations are in Appendix B (Fig. 11).
Sensitivity to scene-target consistency is non-monotonic
We further inspect how the attention-to-target ratio ($\bar{r}_{l}$) varies as a continuous function of scene–object semantic fit. For both COOCo and VISIONS, we restrict analyses to the samples where the model correctly identifies the target. In COOCo, we focus on manipulated images and exclude the original and same-target controls, as these do not involve modifications to the scene semantics; in VISIONS, we use the human-validated object-consistency ratings. We compute the mean attention-to-target ratio for each layer across the selected samples and identify the subset of layers that consistently show higher target focus (top quartile by mean attention). For each image, we then obtain a single attention score by averaging the attention-to-target ratio over this selected subset of layers. Semantic relatedness (COOCo) and object consistency (VISIONS) are normalised to the $[0,1]$ range and treated as continuous predictors. We model their relationship to the aggregated attention-to-target ratio using both linear and quadratic regression functions, and compare their goodness-of-fit via $R^{2}$. We present results for the 0 noise condition in Fig. 8; for results for the noised conditions refer to Appendix B, Fig. 12. At 0 noise, across datasets, attention toward the target exhibits a non-linear dependency on semantic fit. In VISIONS, a linear model revealed a significant negative association between object consistency and attention ($\beta=-0.0858$, $p<.001$), and adding a quadratic term significantly improved fit ($\beta_{2}=0.4395$, $p<.001$; nested model comparison: $F(1,597)=27.75$, $p=1.93\times 10^{-7}$). The resulting curve was convex, with a minimum at $x=0.659$ on the normalized scale, indicating reduced attention for moderately consistent objects and increased attention at higher consistency levels. A similar pattern emerged in COOCo. The linear model again showed a significant negative slope ($\beta=-0.1519$, $p<.001$), and the quadratic term provided a significant improvement ($\beta_{2}=0.2003$, $p<.001$; nested model comparison: $F(1,12051)=19.86$, $p=8\times 10^{-6}$), with the minimum located at $x=0.764$. While the effect is weaker in COOCo, both datasets exhibit reliable convex trends, suggesting that the semantic relatedness between target object and scene modulates attention in a non-monotonic manner rather than through a purely linear decline. This pattern is also present in other conditions for different noise levels and areas.
COOCo dataset
VISIONS dataset
Figure 8: Mean attention-to-target ratio as a function of semantic fit in COOCo (left) and VISIONS (right), aggregated over layers with consistently higher target focus. Points represent individual samples; dashed lines show linear regression fits, and solid red lines show quadratic fits.
@@ 7 Discussion
The RefCLIPScore and Accuracy analyses reveal that models consistently exhibit superior performance when either the original target object or a semantically similar one is present. While it is possible that COCO images were included in models’ training data, the observation that scores are slightly higher for the same target condition than for the original condition suggests that models benefit from small semantic shifts within the same category, potentially because the generated object is more prototypical and thus easier to categorise. Alternatively, inpainting introduces artefacts that may make the target object more salient, thereby drawing additional model attention. This could be mitigated in future work by using more advanced, frontier inpainting models. However, the decrease in performance with reduced object–scene relatedness indicates that models are sensitive to target-context semantic compatibility. Noise manipulations provide additional insight: When targets are weakly related to the scene, reducing contextual information improves performance, whereas for semantically congruent targets, preserving scene context is more beneficial when the target region is degraded.
@@
The output versus scene similarity results for LLaVA-OneVision indicate that as semantic relatedness decreases, the model tends to anchor outputs more to the semantics of the target rather than the broader scene. When target objects are corrupted by noise, correct responses shift towards a stronger reliance on scene context, implying an adaptive strategy that leverages global context to compensate for degraded local features, consistent with Junker and Zarrieß (2024).
@@
In the attention analysis, clear differences related to semantic fit emerge in the Transformer decoder’s attention allocation. Correct categorisation consistently involves higher attention to the target, particularly in middle layers (7–15), suggesting that critical semantic processing occurs there, as already noted in interpretability studies on the LLaVA model family Zhang et al. (2025a, b); Golovanevsky et al. (2025). A compelling observation is the inverse relationship between semantic relatedness and attention allocation: less scene-related targets receive more focused attention. Furthermore, noise conditions modulate attention dynamics significantly. Moderate target-specific noise (0.5 level) triggers a compensatory increase in target-focused attention, aiding correct recognition. Conversely, extreme noise (level 1.0) forces models to shift reliance entirely to context. This dynamic attention behaviour highlights an adaptive mechanism within the model, balancing local target analysis against global context integration depending on image clarity and semantic coherence. Finally, LLaVA-One-Vision’s non-monotonic sensitivity to scene–object fit has a notable parallel in work by Damiano et al. (2024), whose eye-tracking study shows that both highly expected and unexpected objects receive increased fixations (see also Çelikkol et al., 2023). The resulting U-shaped relationship between semantic similarity and overt attention has an analogous trend in our results: moderately consistent targets attract reduced attention, whereas both low- and high-consistency targets receive higher focus. While the underlying mechanisms clearly differ, this similarity indicates that the model’s attention patterns resemble key aspects of human scene processing.